Find a if the point (3,a) is on the line that passes through (-2,7) and (5,-3).
Found 2 solutions by rfer, lwsshak3:
Answer by rfer(16322)
Answer by lwsshak3(11628)

Find a if the point (3,a) is on the line that passes through (-2,7) and (5,-3).
Equation of a line: y=mx+b, m=slope, b-y-intercept
slope=∆y/∆x=7-(-3)/(-2,-5)=(7+3)/-2-5=-10/7
equation: y=-10x/7+b
solving b using given coordinates on the line
7=-10*-2/7+b
b=49/7-20/7=29/7
equation: y=-10x/7+29/7
solving for a: plug in x=3 and solve for y
a=-10*3/7+29/7=-30/7+29/7=-1/7